# jina-clip-v2: Multilingual Multimodal Embeddings for Text and Images

### v1

@@ Abstract
Contrastive Language-Image Pretraining (CLIP) is a highly effective method for aligning images and texts in a shared embedding space. These models are widely used for tasks such as cross-modal information retrieval and multi-modal understanding. However, CLIP models often struggle with text-only tasks, underperforming compared to specialized text models. This performance disparity forces retrieval systems to rely on separate models for text-only and multi-modal tasks. In this work, we build upon our previous model, jina-clip-v1, by introducing a refined framework that utilizes multi-task, multi-stage contrastive learning across multiple languages, coupled with an improved training recipe to enhance text-only retrieval. The resulting model, jina-clip-v2, outperforms its predecessor on text-only and multimodal tasks, while adding multilingual support, better understanding of complex visual documents and efficiency gains thanks to Matryoshka Representation Learning and vector truncation. The model performs comparably to the state-of-the-art in both multilingual-multimodal and multilingual text retrieval benchmarks, addressing the challenge of unifying text-only and multi-modal retrieval systems.
@@ 1 Introduction
Contrastive text-image pre-training has become a widely adopted approach to building robust text-image alignment models. These models are particularly effective for tasks like cross-modal retrieval and zero-shot classification (Radford et al., 2021), demonstrating strong generalization capabilities. Conventional text-image pre-training excels at aligning text and image embeddings but struggles with aligning text-to-text embeddings (Koukounas et al., 2024). This is because cross-modal retrieval training is inadequate for text-only retrieval. The texts available are typically image captions, which are generally short and information-poor, and training procedures don’t use standard techniques for training text embeddings, like hard negatives. (Zhang et al., 2024a)
@@
To overcome these limitations, Koukounas et al. (2024) introduced a multi-task, multi-stage contrastive learning approach to simultaneously align text-text and text-image embeddings. Their method involves three stages: optimizing text-text and text-image embeddings for short image-caption and text-text pairs (stage 1), refining alignment using long text-text pairs and detailed image-caption pairs (stage 2) and further enhancing performance with text triplets containing hard negatives and detailed image-caption pairs (stage 3). The resulting model, jina-clip-v1, achieves strong performance on both the cross-modal CLIP Benchmark111https://github.com/LAION-AI/CLIP_benchmark and text embedding MTEB Benchmark (Muennighoff et al., 2023).
Despite its strengths, this model has several limitations. First, it is an English-only multimodal embedding model, which makes it unsuitable for multilingual document retrieval. Second, jina-clip-v1 struggles with images that contain text, tables, graphs, diagrams, and other complex content (Faysse et al., 2024).
@@
To address these challenges, we propose an enhanced text-image pre-training scheme. Our approach incorporates multilingual text-image and text-text pairs, along with text-image pairs containing text and other difficult-to-process complex visual structures. Additionally, we gradually increase image resolution during pre-training to balance computational efficiency and model performance. The resulting model, jina-clip-v2 depicted in Figure 1, not only achieves competitive performance on cross-modal retrieval benchmarks against state-of-the-art models like NLLB-CLIP-SigLIP (Visheratin, 2023b) but also performs comparably to dedicated multilingual text embedding models like jina-embeddings-v3 (Sturua et al., 2024) on the Retrieval and STS tasks of the multilingual MTEB benchmark (Muennighoff et al., 2023). Moreover, due to the inclusion of visually rich training data and the progressive increase in image resolution, jina-clip-v2 demonstrates significantly improved performance on benchmarks for visually rich document retrieval, compared to jina-clip-v1. To reduce vector storage costs, we employ Matryoshka Representation Learning (Kusupati et al., 2024) to enable dimension truncation on the output vectors with minimal performance degradation.
@@ 2 Related Work
Radford et al. (2021) pioneered the dual encoder architecture trained contrastively on image-text pairs, demonstrating impressive zero-shot performance and laying the groundwork for pre-trained vision-language models. Building upon this foundation, Zhai et al. (2023) propose an alternative and more efficient sigmoid objective, while Sun et al. (2023) optimize the model’s dimensions and scale up both dataset and model sizes, achieving state-of-the-art performance on cross-modal tasks. Koukounas et al. (2024) enhance a CLIP-based retriever with strong text-to-text and cross-modal retrieval capabilities, though fall short in addressing retrieval scenarios that involve candidate pools with heterogeneous modalities, due to the modality gap (Liang et al., 2022). To mitigate the modality bias, Lin et al. (2024) take a different approach by fine-tuning a multimodal LLM with modality-aware hard-negative mining, creating a universal multimodal retriever and maintaining competitive text-to-text retrieval performance across diverse tasks.
@@
Extensive research has also focused on extending CLIP’s capabilities to multilingual contexts. To address the lack of image-caption pairs in languages other than English, Carlsson et al. (2022) apply knowledge distillation to retrain the text encoder using machine translation data. NLLB-CLIP (Visheratin, 2023a) leverages Locked-image Tuning (LiT) (Zhai et al., 2022) alongside the NLLB (Costa-jussà et al., 2022) text encoder, achieving state-of-the-art results in both retrieval and classification tasks.
@@
The importance of good information retrieval in Retrieval-Augmented Generation (RAG) systems has motivated rapid progress in multilingual text embedding models. These models are typically based on either a robust encoder-only architecture, such as XLM-RoBERTa (Conneau et al., 2020) or a decoder-only multilingual large language model, such as Mistral 7B (Jiang et al., 2023). Multilingual E5 (Wang et al., 2024) and BGE-M3 (Chen et al., 2024) both use XLM-RoBERTa as their backbone, leveraging extensive multilingual training with instruction tuning and multi-task learning techniques, respectively. mGTE (Zhang et al., 2024b) achieves comparable performance to BGE-M3 using a smaller transformer architecture, enhanced by RoPE (Su et al., 2023) to extend the context length. Similarly, Sturua et al. (2024) employ RoPE to increase the model’s context length and apply LoRA tuning (Hu et al., 2021) to optimize embeddings for downstream tasks.
3 Training
@@
| Input Specification | 8,192 tokens (max) | 512×512 pixels |
@@
The jina-clip-v2 model uses the dual encoder architecture, introduced in the original CLIP (Radford et al., 2021) model and reused in jina-clip-v1 (Koukounas et al., 2024). One encoder is responsible for encoding texts in a 1024-dimensional embedding space, while a second encoder embeds images in the same embedding space.
The text encoder is initialized with the pre-trained Jina-XLM-RoBERTa model weights. Introduced in jina-embeddings-v3 (Sturua et al., 2024), the Jina-XLM-RoBERTa model is a port of the XLM-RoBERTa (Conneau et al., 2020) weights to a modern encoder-only architecture that adds Flash Attention (Dao et al., 2022), rotary positional embeddings (Su et al., 2023) and LoRA (Hu et al., 2021).
For the image encoder, we have opted for the EVA02 (Fang et al., 2023b) family of ViT models, similarly to jina-clip-v1. We have selected a relatively large pre-trained model, similar in size to the text encoder. This model implementation includes 2D rotary positional embeddings (Su et al., 2023) and a memory-efficient attention implementation based on xFormers (Lefaudeux et al., 2022).
Architectural details for both encoders are presented in Table 1.
3.1 Datasets
Similarly to jina-clip-v1 (Koukounas et al., 2024), we construct four different training datasets. Dataset denotes a dataset of text pairs, denotes a dataset of text samples with hard negatives, a multimodal short-caption dataset and a multimodal long-caption dataset. All datasets are multi-lingual.
The text corpora and are introduced in Sturua et al. (2024) and contain data in 30 languages. The text pair corpus consists of query and target pairs from a diverse collection of text-pair datasets. The triplet text corpus is a high-quality dataset with hard negatives from various sources. Each training item contains one annotated positive and seven negative items. Regarding , we focus on Retrieval and Semantic Textual Similarity (STS) datasets for the scope of this work.
Our multi-modal training data comes from multiple sources. For the English part, we rely on the DFN dataset (Fang et al., 2023a), a collection of 400M image-caption pairs, filtered using data filtering networks. For the rest of the languages, we collect 400M samples from CommonPool (Gadre et al., 2023), using multilingual CLIP to identify and discard low-quality items. All images are resized to . These two large-scale datasets are included in and a small part with captions longer than 256 tokens is held out for .
Inspired by visual QA applications (Faysse et al., 2024), We have diversified our training data with PDFs, scientific graphs, infographics and Wikipedia images. The following datasets are included in both and : DocVQA (Mathew et al., 2021b), TatDQA (Zhu et al., 2022), InfographicsVQA (Mathew et al., 2021a), SciGraphQA (Li & Tajbakhsh, 2023), ArXivQA and ArXivCAP (Li et al., 2024), WIT (Srinivasan et al., 2021) and ViDoRe synthetic training data (Faysse et al., 2024). For the QA datasets, to obtain a caption for an image, we concatenate query and answer for each sample. For the WIT dataset, we use the reference caption in and a concatenation of caption, page title, section tile, page description and section description for .
@@ 3.2 Training Stages
Following the strategy proposed in Koukounas et al. (2024), we employ a multi-task, multi-stage training approach, in order to optimize the model for two tasks simultaneously: text-image matching and text-text matching.
Stage 1 focuses on aligning the multi-modal representations while also improving the text representations of the text encoder. We train on and , with a small context length of 77 and an image resolution of to enable large batch sizes. Towards the end, when performance peaks, we switch to a higher image resolution of using positional embedding interpolation as a warm-up for the next stage.
@@
Appendix table 6 specifies the training details for each step.
@@ 3.3 Loss Functions
Equation 3 formulates the loss for each of the three stages. In each stage, we optimize a joint loss function, i.e. the sum of two loss functions, one operating on text representations and one on multi-modal representations. All stages optimize except Stage 3, which uses hard negatives in the text loss branch and thus calculates .
In Equation 3, denotes a batch of text pairs drawn from the dataset of text pairs , while denotes a batch of text samples with negatives drawn from the hard negatives dataset . Similarly denotes a batch of text-image pairs drawn from the multimodal short-caption dataset , while denotes a batch of text-image pairs sampled from the multimodal long-caption dataset .
@@ 3.4 Matryoshka Representation Learning
In all stages of training, each loss function is evaluated on the full 1024-dimension output vectors of the two encoders as well as on truncated vectors of 64, 128, 256, 512 and 768 dimensions. In other words, each loss component on equation 3 is the sum of the same loss function evaluated at different embedding dimensionalities.
This training technique, called Matryoshka Representation Learning, (Kusupati et al., 2024) trains the model to learn the most important features in the vector sequence. This makes it possible, at inference time, to truncate embedding vectors, thereby reducing storage and computation costs when comparing vectors, at a minor performance penalty.
@@ 4 Evaluation
We have evaluated the performance of jina-clip-v2 on a selection of benchmarks: English and multilingual cross-modal retrieval benchmarks in Section 4.1, common text embedding tasks from the MTEB benchmark suite in Section 4.2, visual document retrieval tasks in Section 4.3, and an ablation study of jina-clip-v2’s matryoshka representations in Section 4.4.
Detailed evaluation results are presented in the appendices. Detailed CLIP Benchmark results are given in Appendix tables 7, 8 and 9, MTEB results are in in Appendix 10. ViDoRe benchmark results are in Appendix 11. A full MRL evaluation can be found in tables 12, 13, 14, 15 in the appendices.
4.1 Cross-modal Evaluation
| Benchmark | CLIP Benchmark - EN | Crossmodal-3600 | Multilingual MS COCO | |||
| Task Type | Zero-Shot Retrieval | |||||
| Model - Metric | T-I r@5 | I-T r@5 | T-I r@5 | I-T r@5 | T-I r@5 | I-T r@5 |
| nllb-siglip-large | 81.54 | 88.15 | 82.07 | 80.16 | 87.60 | 85.37 |
| nllb-siglip-base | 79.57 | 86.38 | 79.29 | 76.56 | 86.23 | 84.87 |
| jina-clip-v1 | 77.75 | 87.65 | - | - | - | - |
| jina-clip-v2stage 1 | 73.87 | 86.61 | 73.51 | 79.64 | 80.66 | 83.02 |
| jina-clip-v2stage 2 | 79.81 | 89.57 | 84.13 | 84.12 | 86.11 | 86.45 |
| jina-clip-v2 | 79.09 | 89.73 | 81.43 | 83.23 | 84.87 | 86.03 |

### v2

@@ Abstract
Contrastive Language-Image Pretraining (CLIP) has been widely used for crossmodal information retrieval and multimodal understanding tasks. However, CLIP models are mainly optimized for crossmodal vision-language tasks and underperform in single-mode text tasks. Moreover, these models are often trained on English datasets and therefore lack multilingual understanding. Additionally, from a visual understanding perspective, previous CLIP-based models exhibit insufficient understanding of visually rich documents. In this work, we propose jina-clip-v2, a contrastive vision-language model trained on text pairs, triplets and image-text pairs via a multi-task and multi-stage contrastive learning paradigm in order to support both text-only and crossmodal tasks. We employ a multilingual text encoder and expand the training dataset to include multilingual texts from 29 non-English languages, including Hindi, Chinese, German, French, and others, as well as images of visually-rich documents. We evaluate the model’s performance and show that jina-clip-v2 achieves notable improvements over state-of-the-art CLIP-based models in zero-shot text-only retrieval, semantic textual similarity, and crossmodal retrieval tasks in both English and multilingual settings. jina-clip-v2 also provides for flexibility in embedding dimensionality, enabling users to select the granularity of the representations. jina-clip-v2 is publicly available at https://huggingface.co/jinaai/jina-clip-v2.
@@ 1 Introduction
Contrastive text-image pretraining is a well-known architecture and training framework for building robust text-image alignment models (Radford et al., 2021; Zhai et al., 2023; Sun et al., 2023). These models are particularly effective for tasks like crossmodal retrieval and zero-shot classification, demonstrating strong generalization capabilities. However, although CLIP models produce general-purpose embeddings for both texts and images, they are woefully inadequate as a source of text embeddings (Koukounas et al., 2024). We believe this is due to the training regimen of CLIP models, which are typically trained with short and information-poor image captions (Zhang et al., 2024a) and omit standard techniques for finetuning meaningful text embeddings, e.g. training using hard negatives (Gillick et al., 2019; Qu et al., 2021).
@@
To overcome these limitations, Koukounas et al. (2024) introduced a multi-task, multi-stage contrastive learning approach to simultaneously align text-text and text-image embeddings. This method involved three stages: optimizing text-text and text-image embeddings for short image-caption and text-text pairs (stage 1), refining alignment using long text-text pairs and detailed image-caption pairs (stage 2), and further enhancing performance with text triplets containing hard negatives and detailed image-caption pairs (stage 3). The resulting model, jina-clip-v1, achieves strong performance on both the crossmodal CLIP Benchmark111https://github.com/LAION-AI/CLIP_benchmark and the text embedding MTEB Benchmark (Muennighoff et al., 2023). Despite its strengths, jina-clip-v1 (Koukounas et al., 2024) has several limitations. First, it is an English-only model, which makes it unsuitable for multilingual document retrieval. Second, jina-clip-v1 struggles with visually-rich images, e.g. those containing text, tables, graphs, and diagrams (Faysse et al., 2024). Moreover, jina-clip-v1 only supports fixed length embeddings, which can be wasteful when embedding vectors are much larger than required for a task (Kusupati et al., 2024).
To address these challenges, we propose several enhancements to the pretraining scheme. For multilingual support, our approach uses a multilingual language model (Conneau et al., 2020) to initialize the text encoder, and incorporates multilingual text-image and text-text pairs during training. To improve performance on visually complex documents, we gradually increase image resolution during pretraining and use multimodal training pairs with complex visual structures. Finally, to enable users to select the granularity of their representations, we employ Matryoshka Representation Learning (Kusupati et al., 2024) and enable dimensional truncation on the output vectors with minimal performance degradation.
@@
The resulting model, jina-clip-v2 depicted in Figure 1, not only achieves competitive performance on crossmodal retrieval benchmarks against multilingual state-of-the-art models like NLLB-CLIP (Visheratin, 2023b) but also performs comparably to dedicated multilingual text embedding models like jina-embeddings-v3 (Sturua et al., 2024) on the retrieval and semantic textual similarity (STS) tasks of the MTEB benchmark (Muennighoff et al., 2023). Moreover, due to the inclusion of visually rich training data and the progressive increase in image resolution, jina-clip-v2 demonstrates significantly improved performance on ViDoRe, a benchmark for visually rich document retrieval (Faysse et al., 2024), compared to jina-clip-v1. In summary, our contributions are the following:
1.
Support for multiple languages. jina-clip-v2 achieves up to 67% higher scores on multilingual crossmodal performance, up to 60% on multilingual Retrieval performance and up to 43% on multilingual STS, compared jina-clip-v1.
2.
Improved performance on visual document retrieval. jina-clip-v2’s scores are 35% higher on the ViDoRe benchmark (Faysse et al., 2024) compared to jina-clip-v1.
3.
Flexible embedding dimensionality. Truncating jina-clip-v2 embeddings by up to 75% (from 1,024 dimensions down to 256) reduces performance by <1%.
@@ 2 Related Work
Previous work (Jia et al., 2021; Li et al., 2021; Radford et al., 2021; Zhang et al., 2022) pioneered dual-encoder architectures trained contrastively on image-text pairs, demonstrating impressive zero-shot performance and laying the groundwork for pretrained vision-language models. Building on this foundation, Zhai et al. (2023) proposed an alternative and more efficient sigmoid objective, while Sun et al. (2023) optimized the model’s dimensions and scaled up both dataset and model sizes, achieving state-of-the-art performance on crossmodal tasks. Koukounas et al. (2024) enhanced a CLIP-based retriever with strong text-to-text and crossmodal retrieval capabilities, however, their work fails to address retrieval scenarios involving candidate pools with heterogeneous modalities, due to the modality gap (Liang et al., 2022; Eslami & de Melo, 2024) and the modality bias (Lin et al., 2024). To mitigate the modality bias, Lin et al. (2024) finetuned a multimodal LLM with modality-aware hard-negative mining, creating a universal multimodal retriever and maintaining competitive text-to-text retrieval performance across diverse tasks. In this work, we build on the training strategy proposed by Koukounas et al. (2024) and train our model using a text-to-text as well as a text-to-image contrastive InfoNCE (Van den Oord et al., 2018) loss.
@@
Extensive research has also focused on extending CLIP to languages beyond English. To address the lack of image-caption pairs in languages other than English, Carlsson et al. (2022) applied knowledge distillation to retrain the text encoder using machine-translated data. NLLB-CLIP (Visheratin, 2023a) leverages Locked-image Tuning (LiT) (Zhai et al., 2022) alongside the multilingual NLLB (Costa-jussà et al., 2022) text encoder, achieving state-of-the-art results in both retrieval and classification tasks.
@@
Multilingual text embedding models are typically based on either an encoder-only architecture, such as XLM-RoBERTa (Conneau et al., 2020), or a decoder-only multilingual large language model, such as Mistral 7B (Jiang et al., 2023). Multilingual E5 (Wang et al., 2024) and BGE-M3 (Chen et al., 2024) both use XLM-RoBERTa as their backbone, the first leveraging extensive multilingual training with instruction tuning, and the second employing multi-task learning techniques. mGTE (Zhang et al., 2024b) achieves comparable performance to BGE-M3 using a smaller transformer architecture, enhanced by RoPE (Su et al., 2023) to extend the input context length. Similarly, Sturua et al. (2024) employ RoPE to increase the model’s context length and apply LoRA tuning (Hu et al., 2021) to optimize embeddings for downstream tasks.
Finally, Kusupati et al. (2024) propose a training technique, which they call Matryoshka Representation Learning, that enables embedding models to learn coarse-to-fine representations. These representations can be truncated during inference to match the requirements of downstream tasks, cutting down on potentially redundant vector dimensions.
3 jina-clip-v2
The jina-clip-v2 model uses the dual encoder architecture, introduced in the original CLIP (Radford et al., 2021) model and reused in jina-clip-v1 (Koukounas et al., 2024).
The text encoder is initialized with pretrained Jina-XLM-RoBERTa model weights. Introduced in jina-embeddings-v3 (Sturua et al., 2024), the Jina-XLM-RoBERTa model is a port of the multilingual XLM-RoBERTa (Conneau et al., 2020) model to a modern encoder-only architecture with Flash Attention (Dao et al., 2022), rotary positional embeddings (Su et al., 2023) and support for low-rank adaptation (Hu et al., 2021).
Like jina-clip-v1, the image encoder is a pretrained EVA02 model (Fang et al., 2023b). We selected the L/14 pretrained model variant, which is similar in number of parameters to the text encoder. This model implementation includes 2D rotary positional embeddings and a memory-efficient attention implementation based on xFormers (Lefaudeux et al., 2022). Table 1 presents architectural details for both encoders.
@@
| Input Specification | 8,192 tokens (max) | resolution |
@@
3.1 Training Data
Similarly to jina-clip-v1 (Koukounas et al., 2024), we constructed four training datasets used in different stages of the training process. Dataset denotes a dataset of text pairs, a dataset of text samples with hard negatives, a multimodal short-caption dataset and a multimodal long-caption dataset.
Both and contain data in 30 languages drawn from various existing datasets for text information retrieval and semantic similarity, and were introduced as training data for jina-embeddings-v3 (Sturua et al., 2024). consists of pairs of texts drawn from a diverse collection of datasets, and is a high-quality dataset with hard negatives from various sources, where each sample contains one annotated positive and seven negative items. Sturua et al. (2024) provide detailed information about these datasets and how they were curated.
Our multimodal training datasets draw on multiple sources. We randomly sampled 400M image-caption pairs from the 2B in the DFN dataset (Fang et al., 2023a) to obtain a collections of image-text pairs in English. To this, we added non-English image-text pairs sampled from CommonPool (Gadre et al., 2023). We filtered this data by language, aspect ratio, and by removing any caption with less than five words. We used multilingual SigLIP (Zhai et al., 2023) to rank image-text pairs based on cosine similarity and kept the top scoring 50%. The result was an additional 400M image-text pairs. All images were resized to . These two large-scale datasets are included in and a small part with captions longer than 256 tokens is held out for .
In order to overcome the limitations of visual document understanding, we diversify our training data with PDFs, scientific graphs, infographics and Wikipedia images. Specifically, the short caption image-text dataset and the long-caption include the following training datasets: DocVQA (Mathew et al., 2021b), TatDQA (Zhu et al., 2022), InfographicsVQA (Mathew et al., 2021a), SciGraphQA (Li & Tajbakhsh, 2023), ArXivQA and ArXivCAP (Li et al., 2024), WIT (Srinivasan et al., 2021) and ViDoRe synthetic training data (Faysse et al., 2024). For the QA datasets, we concatenated the query and answer for each sample to construct a matching text in lieu of a caption. For the WIT dataset, we use the reference caption as corresponding text in and a concatenation of reference caption, page title, section title, page description and section description as matching text for .
@@ 3.2 Training Stages
Inspired by Koukounas et al. (2024), we employ a multi-task, multi-stage training approach, in order to optimize the model for two tasks simultaneously: text-image matching and text-text matching.
Stage 1 focuses on aligning the multimodal representations while also improving the text representations of the text encoder. We train on and , with a small context length of 77 and an image resolution of to enable large batch sizes of . This is the most computationally expensive stage in our training pipeline, requiring approximately 10 days on 8 NVIDIA H100 GPUs to converge. Towards the end, when performance peaks, we switch to a higher image resolution of using positional embedding interpolation as a warm-up for the next stage.
@@ 3.3 Loss Functions
Equation 3 defines the loss for each of the three stages. In each stage, we optimize a joint loss function, i.e. the sum of two loss functions, one operating on text representations and one on multimodal representations. In equation 3, denotes a batch of samples drawn from the dataset , where . All stages optimize except Stage 3, which uses hard negatives in the text loss branch and thus calculates .
@@ 3.4 Matryoshka Representation Learning
In every training stage, the loss component is recalculated at different dimensionalities: Loss is computed for the full 1024-dim output vector, and for truncated subsets of 64, 128, 256, 512 and 768 dimensions. Model weight adjustment proceeds on the basis of all those losses combined. This training technique, called Matryoshka Representation Learning (Kusupati et al., 2024), induces the model to encode its embeddings by increasing granularity. This makes it possible, at inference time, to truncate embedding vectors with a minor performance penalty.
@@ 4 Evaluation
We evaluate the performance of jina-clip-v2 on a selection of benchmarks: English and multilingual crossmodal retrieval benchmarks in Section 4.1, text embedding tasks from the MTEB benchmark suite in Section 4.2 and visual document retrieval tasks in Section 4.3. Finally, we present an ablation study of jina-clip-v2’s Matryoshka representations in Section 4.4. Detailed evaluation results are presented in Section A tables 7 to 26.
4.1 Crossmodal Retrieval Evaluation
| Benchmark | CLIP Benchmark | Crossmodal-3600 | XTD10 | ||||
| Language | English | Multilingual | Multilingual | ||||
| Task Type | Zero-Shot Retrieval | ||||||
| Model | #Parameters | T-I r@5 | I-T r@5 | T-I r@5 | I-T r@5 | T-I r@5 | I-T r@5 |
| jina-clip-v1 | 223M | 77.75 | 87.65 | 16.93 | 19.82 | 31.01 | 36.89 |
| nllb-siglip-base | 507M | 79.57 | 86.38 | 79.29 | 76.56 | 86.23 | 84.87 |
| nllb-siglip-large | 1.2B | 81.54 | 88.15 | 82.07 | 80.16 | 87.60 | 85.37 |
| jina-clip-v2stage 1 | 865M | 73.87 | 86.61 | 73.51 | 79.64 | 80.66 | 83.02 |
| jina-clip-v2stage 2 | 865M | 79.81 | 89.57 | 84.13 | 84.12 | 86.11 | 86.45 |
| jina-clip-v2 | 865M | 79.09 | 89.73 | 81.43 | 83.23 | 84.87 | 86.03 |